OpenStudy (anonymous):

9x-15y=3 i need help finding x and y?

OpenStudy (anonymous):

in terms of what?

OpenStudy (anonymous):

im not sure it just asks to find x and y

OpenStudy (anonymous):

any x and y? you know the equation you wrote is that of a straight line,and there are an infinite number of points on a straight line. if you need some of these points just plug in a value of x and find the corresponding value of y ex: if x=0 then -15y=3 and y=-3/15=-1/5 etc.

OpenStudy (anonymous):

9x-15y=3 let us just give values for x and substitute it in the given equation and find the value of yand so on!!!